An algebric equation which is reducible to qudratic equation:
X^6 - 9X^3 + 8 =0

Answer by RAY100(1637)

x^6 -9x^3 +8 =0
.
let y=x^3
.
y^2 -9y +8 =0
.
(y-8)(y-1) =0
.
y=8,,1
.
but,,,y=x^3
.
x^3 =8
.
x=2
.
x^3 =1
.
x=1
.